IF A TRIANGLE HAS A HEIGHT OF 5 CENTIMETERS AND A BASE OF 3 CENTIMETERS THEN THE AREA EQUALS ?
Answer by lawdej(58)

area=1/2base x height
=1/2 x5cmx3cm
=15/2cm^2
=7.5 cm^2